OpenStudy (anonymous):

given SEC= 8/3, find the 6 trigonometric ratios for beta

OpenStudy (anonymous1):

If the secant is 8/3, the cosine is the inverse of the secant: COS = 1/SEC COS = 3/8 Is this what you're asking?

OpenStudy (anonymous):

\[\sec(\beta)=\frac{8}{3}\] \[\cos(\beta)=\frac{3}{8}\] the others depend on what quadrant you are in

OpenStudy (anonymous):

we can pretend we are in quadrant one and solve those

OpenStudy (anonymous1):

For the sine you can use: \[\sin^2\beta+\cos^2\beta=1\] But, as satellite73 said, for the sign of the sine the quadrant is important.

OpenStudy (anonymous1):

Since the cosine is positive, the quadrants are limited 1 and 4. So, we can do what satellite73 said and pretend we are in quadrant 1.

OpenStudy (anonymous):

that's actully wrong.

OpenStudy (anonymous1):

How so?